A wave on a string has an equation

x = (0.15 metres) cos (25t-5.0x) .

Assuming all values are in the appropriate SI units, what is the speed of this wave?

Select one:

a. 2.0 m/s

b. 5.0 m/s

c. 7.0 m/s

d. 18 m/s

e. 35 m/s
Answer #1

here,

x = (0.15 m) * cos(25 t - 5x)

compairing the equation with

x = A * cos(w * t - K * x)

w = 25 rad , k = 5 m^-1

the wavelength , lamda = 2*pi /K

lamda = 2*pi /5 = 1.256 m

the frequency , f = w /2*pi

f = 25 /(2*pi) = 3.98 Hz

the speed of this wave , v = f * lamda

v = 3.98 * 1.256 m/s = 5 m/s

the speed of this wave is b) 5 m/s